4. The inlet tube that supplies air pressure to operate a hydraulic jack has a 30mm radius. The output piston has a 64mm radius. What air pressure should be used to lift with the crane a 2300kg mass?
1
Expert's answer
2018-09-07T16:55:08-0400
We can find the air pressure that should be used to lift with the crane from the hydraulic press formula:
P_1=F_1/A_1 =F_2/A_2 =P_2,
here, P_1=P_2 is the air pressure that should be used to lift with the crane, F_1 is the force that act on a small area of the inlet tube, F_2 is the force that act on the output piston, A_1 is the cross-sectional area of the inlet tube, A_2 is the cross-sectional area of the output piston.
Then, we get:
P_1=F_2/A_2 =mg/(πr^2 )=(2300 kg∙9.8 m/s^2 )/(π∙(0.064 m)^2 )=1.75∙〖10〗^6 Pa.